OpenStudy (anonymous):

What is the solution of 12/b = 36/51?

OpenStudy (anonymous):

cross multiply.

OpenStudy (anonymous):

36*b = 12*51

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

b = (12*51)/(36)

OpenStudy (anonymous):

hopefully this helps and makes sense....

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):